Safawi kicked a ball with an initial speed of 20 ms-1 at 30º angle. If the ball experienced a constant vertical acceleration of -9.81 ms-2, calculate the ball’s maximum height and distance. Explain how soccer player could apply the force of summation theory to optimize his kicks performance.  


1
Expert's answer
2022-02-08T15:10:16-0500

Explanation & Calculations


  • Apply "\uparrow\,v^2 = u^2+2as" to calculate the max height.

"\qquad\qquad
\begin{aligned}
\small 0^2&=\small (20\sin30\,ms^{-1})^2+2(-9.81\,ms^{-1}\times h)\\
\small h&=\small 5.1\,m
\end{aligned}"


  • Apply "\small s = \frac{u^2}{2g}\sin 2\theta" horizontally to calculate the distance.


  • The player moves the dominant leg bent at first and extends it just before hitting the ball.
  • For this to happen good, the leg should be bent sufficiently.
  • The bent leg is easier to move around the hip than the extended leg.


  • The opposite hand on the other side needs to be kept fully extended while the leg is bent and hatched while the leg is fully extended.